A monopolist with constant average and marginal cost equal to 10 (AC = MC = 10) faces demand Q = 100 - 2P, implying that its marginal revenue is MR = 100 - 4Q.  Its profit maximizing quantity is

Group of answer choices

22.5

45

90

80
Answer #1

A monopolist produces at the point where MR= MC.

Equating the given values: 100-4Q = 10

100-10= 4Q

90 = 4Q

90/4 = Q

22.5 = Q

The correct answer is: (i) 22.5